# Cosmology from one galaxy in voids?

### v1

@@ 3.2 Core Density Cut and Radius Cut
To further reduce the number of spurious voids in the void catalog, in this Section, we remove voids also based on the radius cut (besides the core density cut). Similar to Section 3.1, Figure 3 shows the results of non-void galaxies, random galaxies and void galaxies to infer the values of $\Omega_{\mathrm{m}}$. The values of $R^{2}$ and RMSE have a stronger improvement after this additional purity cut on void radius is applied, showing an even higher constraining power of void galaxies. In addition, in this case, we see that the $\epsilon$ value also improves when considering more isolated galaxies (bottom row, right plot). Therefore, when both criteria for purity are considered (core density cut and void radius), $\epsilon$ improves both when we select void galaxies instead of non-void or randomly selected galaxies (moving from the left and center panels to the right panels) and when we increase the Voronoi cell cut-off volume (from up to bottom panel for void galaxies). By comparing Figure 2 and Figure 3, we conjecture that, since the radius cut removes outlier
galaxies with large error bars, it may be responsible for this improvement. If this is the case, the smaller error bars would be the reason for a smaller $\epsilon$ value. As in the case with only the core density cut, the strongest improvement in constraints is obtained when accounting for selections on both the large-scale and local environments (right column, bottom plot). Therefore, the results of this Section confirm that when using a galaxy from a cosmic void we obtain more stringent constraints.
@@ 3.3 Test on the mean values of predictions with galaxies from the same simulation
In Villaescusa-Navarro et al. (2022), the authors show the average prediction of galaxies in the same simulation. Here, to investigate the constraining capabilities on average of the different kinds of galaxies separately we show the average predictions of void, non-void and randomly selected galaxies from the same simulation (see the top row in Figure 4). We note that we perform this binning only for the case in which both purity cuts are applied (radius and core density). In this case, we see a milder improvement when using void galaxies, with the value of $\epsilon$ being slightly boosted. If, however, we zoom in to the lower value of $\Omega_{\mathrm{m}}$ as shown in the second row of Figure 4, we notice that the error bars on the lower values of $\Omega_{\mathrm{m}}$ are much smaller when considering void galaxies (seeing an improvement in the bottom row from left to right), corresponding to a further improvement of the $\epsilon$ values. This shows that void galaxies seem to particularly provide better constraints for the lower values of $\Omega_{\mathrm{m}}$, roughly between 0.1 and 0.2. Yet, we notice that there are some biases for the lower values of $\Omega_{\mathrm{m}}$, as well as closer to the higher extreme values of $\Omega_{\mathrm{m}}$.
Furthermore, in Figure 5 we plot the error bars of the predictions as a function of the Voronoi cell volume of the galaxies (considering galaxies within 75% of the void radius from void centers as above). A correlation is present: we notice that the error bars of the predictions (x-axis) are smaller for galaxies with larger Voronoi cell volume (y-axis)—indicating that more isolated void galaxies are more constraining. This provides additional evidence that galaxies in underdense regions are better for predicting $\Omega_{\mathrm{m}}$.
@@ 4 Conclusion
In this work, we have shown that the correlation between $\Omega_{\rm m}$ and galaxy properties exhibit less scatter for void galaxies than for non-voids galaxies. This is an important and new result that may shed light on the impact of baryonic processes on galaxy formation and evolution as a function of the cosmic environment. Recent works showing that void galaxies have distinct characteristics in the IllustrisTNG simulation (Rodríguez-Medrano et al., 2023a; Curtis et al., 2024) could hint to an explanation for our results. We also show that the improvement obtained when using void galaxies is more significant when a more stringent selection is considered, robustly selecting according to both the large-scale and local environments. This improvement, seen when more conservatively selecting void galaxies, seems to confirm the physical expectation that void galaxies provide a cleaner environment to extract cosmological information. Additionally, we notice that the constraining power of void galaxies seems higher in the case of lower values of $\Omega_{\mathrm{m}}$, a result that requires further investigation and paves the way to future studies on the topic.
Our results imply that, if we consider void galaxies, we can obtain similar accuracy and precision for cosmological inference as when using non-void or randomly selected galaxies, but with much less data. It also shows that, for an equal number of galaxies selected, the constraining power obtained by galaxies in voids is stronger than the one obtained when using random galaxies. The implications of our results therefore favor the interpretation that void galaxies provide a cleaner and more effective environment for extracting cosmological information. These results should be further tested on different simulations, such as ASTRID and SIMBA. Looking ahead to applications to data, it would be interesting to investigate whether these results are expected to hold with observational data from e.g. photometric surveys (see Hahn et al., 2023). We notice that, with respect to observation strategies, there is a trade-off between using all available galaxies versus restricting to void galaxies: while void galaxies yield better constraints, they are scarcer. For better constraints, the strategy could directly aim at targeting void galaxies. Finally, our results may, in the longer term, motivate future observation campaigns to observe sky patches in underdense regions in order to extract cosmological information.
@@ 5 Acknowledgments
We are especially grateful to Francisco Villaescusa-Navarro, whose invaluable help and support were pivotal to the completion of this paper. We also thank Wenhan Zhou, Tri Nguyen, Eiichiro Komatsu, Leander Thiele, ChangHoon Hahn, and Sohaib Bhatti for valuable discussions. BW and AP acknowledge support from the Simons Foundation to the Center for Computational Astrophysics at the Flatiron Institute. AP acknowledges support from the European Research Council (ERC) under the European Union’s Horizon programme (COSMOBEST ERC funded project, grant agreement 101078174).

### v2

@@ 3.2 Core Density Cut and Radius Cut
To further reduce the number of spurious voids in the void catalog, in this Section, we remove voids also based on the radius cut (besides the core density cut). Similar to Section 3.1, Figure 3 shows the results of non-void galaxies, random galaxies and void galaxies to infer the values of $\Omega_{\mathrm{m}}$. The values of $R^{2}$ and RMSE have a stronger improvement ($R^{2}\sim 8\%$ and RMSE $\sim 40\%$ for comparing random galaxies and void galaxies in the bottom row) after this additional purity cut on void radius is applied, showing an even higher constraining power of void galaxies. In addition, in this case, we see that the $\epsilon$ value also improves when considering more isolated galaxies (bottom row, right plot). Therefore, when both criteria for purity are considered (core density cut and void radius), $\epsilon$ improves both when we select void galaxies instead of non-void or randomly selected galaxies (moving from the left and center panels to the right panels) and when we increase the Voronoi cell cut-off volume (from up to bottom panel for void galaxies). By comparing Figure 2 and Figure 3, we conjecture that, since the radius cut removes outlier galaxies with large error bars, it may be responsible for this improvement. If this is the case, the smaller error bars would be the reason for a smaller $\epsilon$ value. As in the case with only the core density cut, the strongest improvement in constraints is obtained when accounting for selections on both the large-scale and local environments (right column, bottom plot). Therefore, the results of this Section confirm that when using a galaxy from a cosmic void we obtain more stringent constraints.
@@ 3.3 Test on the mean values of predictions with galaxies from the same simulation
In Villaescusa-Navarro et al. (2022), the authors show the average prediction of galaxies in the same simulation. Here, to investigate the constraining capabilities on average of the different kinds of galaxies separately (but irrespective of variations of the model of the simulation), we show the average predictions of void, non-void and randomly selected galaxies from the same simulation, that is a simulation where the true value of $\Omega_{\mathrm{m}}$ is the same, as well as other parameters (see the top row in Figure 4). We note that we perform this binning only for the case in which both purity cuts are applied (radius and core density). In this case, we see a milder improvement when using void galaxies, with the value of $\epsilon$ being slightly boosted. If, however, we zoom in to the lower value of $\Omega_{\mathrm{m}}$ as shown in the second row of Figure 4, we notice that the error bars on the lower values of $\Omega_{\mathrm{m}}$ are much smaller when considering void galaxies (seeing an improvement in the bottom row from left to right), corresponding to a further improvement of the $\epsilon$ values. This shows that void galaxies seem to particularly provide better constraints for the lower values of $\Omega_{\mathrm{m}}$, roughly between 0.1 and 0.2. Yet, we notice that there are some biases for the lower values of $\Omega_{\mathrm{m}}$, as well as closer to the higher extreme values of $\Omega_{\mathrm{m}}$.
Furthermore, in Figure 5 we plot the error bars of the predictions as a function of the Voronoi cell volume of the galaxies (considering galaxies within 75% of the void radius from void centers as above). A correlation is present: we notice that the error bars of the predictions (x-axis) are smaller for galaxies with larger Voronoi cell volume (y-axis)—indicating that more isolated void galaxies are more constraining. The scale at which a void galaxy is isolated (from a local environment perspective) corresponds to Voronoi cell volumes of 40 $(\mathrm{Mpc}/h)^{3}$, that is to a scale of $\sim 5\mathrm{Mpc}$, considering $h\approx 0.7$. This provides additional evidence that galaxies in underdense regions are better for predicting $\Omega_{\mathrm{m}}$.
@@ 4 Conclusion
In this work, we have shown that the correlation between $\Omega_{\rm m}$ and galaxy properties exhibit less scatter for void galaxies than for non-voids galaxies. This is an important and new result that may shed light on the impact of baryonic processes on galaxy formation and evolution as a function of the cosmic environment. Recent works showing that void galaxies have distinct characteristics in the IllustrisTNG simulation (Rodríguez-Medrano et al., 2023; Curtis et al., 2024) could hint to an explanation for our results. We also show that the improvement obtained when using void galaxies is more significant when a more stringent selection is considered, robustly selecting according to both the large-scale and local environments. This improvement, seen when more conservatively selecting void galaxies, seems to confirm the physical expectation that void galaxies provide a cleaner environment to extract cosmological information. Additionally, we notice that the constraining power of void galaxies seems higher in the case of lower values of $\Omega_{\mathrm{m}}$, a result that requires further investigation and paves the way to future studies on the topic.
Our results imply that, if we consider void galaxies, we can obtain similar accuracy and precision for cosmological inference as when using non-void or randomly selected galaxies, but with much less data. It also shows that, for an equal number of galaxies selected, the constraining power obtained by galaxies in voids is stronger than the one obtained when using random galaxies. The implications of our results therefore favor the interpretation that void galaxies provide a cleaner and more effective environment for extracting cosmological information. These results should be further tested on different simulations, such as ASTRID and SIMBA.
Looking ahead to applications to data, it would be interesting to investigate whether these results are expected to hold with observational data from e.g. photometric surveys (see Hahn et al., 2023). Our results can serve as an indication to observers wishing to robustly select void galaxies by taking into account both the large-scale void structure and the local environment. Additionally, we notice that, with respect to observation strategies, there is a trade-off between using all available galaxies versus restricting to void galaxies: while void galaxies yield better constraints, they are scarcer. For more stringent constraints, the strategy could directly aim at targeting void galaxies. Finally, our results may, in the longer term, motivate observation campaigns such as CAVITY (Pérez et al., 2024), DIVE (de los Reyes et al., 2023) and Nearby Voids (Pustilnik et al., 2019) to observe sky patches in underdense regions, but specifically designed to extract cosmological constraints.
@@ 5 Acknowledgments
We are especially grateful to Francisco Villaescusa-Navarro, whose invaluable help and support were pivotal to the completion of this paper. We also thank Wenhan Zhou, Tri Nguyen, Eiichiro Komatsu, Leander Thiele, ChangHoon Hahn, and Sohaib Bhatti for valuable discussions. We are grateful to the anonymous referee for their constructive report that considerably improved this paper. BW and AP acknowledge support from the Simons Foundation to the Center for Computational Astrophysics at the Flatiron Institute. AP acknowledges support from the European Research Council (ERC) under the European Union’s Horizon programme (COSMOBEST ERC funded project, grant agreement 101078174).